OpenStudy (anonymous):

Mathematical Induction Help: please help I am really lost and I need step by step to understand. :)

OpenStudy (anonymous):

4−6+5−7+6−8+...+4n(4n+2)=4n(4n+1)(8n+7)6\[4−6+5−7+6−8+...+4n(4n+2)=\frac{ 4(4n+1)(8n+7) }{ 6}\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@shreehari499